OpenStudy (anonymous):

The sum of Lexie’s age and Melissa’s age is 48. Lexie’s age is 6 years less than twice Melissa’s age. Find Lexie’s and Melissa’s ages. Lexie is 32 years old, Melissa is 16 years old. Lexie is 42 years old, Melissa is 6 years old. Lexie is 26 years old, JMelissa is 22 years old. Lexie is 30 years old, Melissa is 18 years old.

OpenStudy (angela210793):

the logic is the same as at the problem with garden...

OpenStudy (anonymous):

L + M = 48 ...1 L = 2M -6 ...2 replace 2 in 1 then 2M - 6 + M = 48 3M -6 =48 3M = 54 M = 18 by equation 2. L = 2M - 6 , we know that the age of M = 18 so, L = 2(18) -6 L = 30

OpenStudy (anonymous):

thanks boo (: